OpenStudy (anonymous):

how can i create a polynomial my own third degree polynomial that when divided by x + 2 has a remainder of –4?

OpenStudy (jamesj):

Suppose P(x) is your polynomial. If when divided by (x+2) it has a remainder of -4, then that means we can write P(x) = (x+2)Q(x) - 4 for some other polynomial Q(x). Hence to answer your question, choose any polynomial Q(x) such that P(x) is third order. That is, make Q(x) an arbitrary second order polynomial.

OpenStudy (anonymous):

what would the equation look like when written out?

OpenStudy (jamesj):

Well, after you've chosen a Q(x), you can expand it and see. What's the simplest second order polynomial Q(x) you can think of?

OpenStudy (anonymous):

i have no idea

OpenStudy (jamesj):

What is an example of a second order polynomial?

OpenStudy (jamesj):

What's the definition of a nth order polynomial?

OpenStudy (jamesj):

A polynomial is a finite sum of monomials, such as x, 2x, -5x, x^3, -17x^4, etc. The order of a polynomial is the highest power of x in the monomials. For example x, -2x, 17x, 5x + 3 are all examples of first order polynomials. x^2, -3x^2, x^2 + x, x^2 + 5, x^2 - 18x + 25959 are all examples of second order polynomials x^3 is a third order polynomial, as is (x+1)^3 = x^3 + 3x^2 + 3x + 1 Make sense?

OpenStudy (anonymous):

-x^3-3x^2-2x this is what i got !

OpenStudy (anonymous):